OpenStudy (anonymous):

My answer is 6. Please help me check. Sam hiked up a mountain at a rate of 2.5 mph and returned along the same trail at a rate of 3mph. His total hiking time was 11h. How long was the hiking trail?

OpenStudy (anonymous):

\[t = \frac ds\] \[11 = \frac{d}{2.5} + \frac{d}{3}\] \[11 = \frac{6d}{15} + \frac{5d}{15}\] \[11 = \frac{11d}{15}\] \[161 = 11d\] \[15 = d\]

OpenStudy (anonymous):

Hold on. Don't mind the last few steps. It should be 165 = 11d d = 15

OpenStudy (anonymous):

Sorry for late reply. Where does 165 comes from?